# More is better: Strong constraints on the stellar properties of LEGA-C galaxies with Prospector

### v1

@@ 4.3 The color-color diagram
The various, complex physical processes that take place in galaxies give rise to the differences observed at UV and optical wavelengths. Interpreting the correlations between optical colors of galaxies is a challenging task. The combined effects of dust attenuation, star formation, and stellar evolution are very difficult to disentangle, leading to many degeneracies between these parameters. A simple yet powerful method to break these degeneracies is to combine an optical and an optical–NIR color. The restframe color-color diagram has proven to be quite successful in that regard (Labbé et al. 2005; Wuyts et al. 2007; Williams et al. 2009). A diagram not only can break the aforementioned degeneracies, but also can classify galaxies into quiescent and star forming. The diagram is commonly used as a quiescence diagnostic in both observed galaxy samples (e.g. Whitaker et al. 2013; Straatman et al. 2014, 2016; Barro et al. 2014; Papovich et al. 2015, 2018; Fang et al. 2018; Tan et al. 2022; Miller et al. 2022; Antwi-Danso et al. 2023; Valentino et al. 2023) and in and cosmological hydrodynamical simulations (Davé et al. 2017; Trayford et al. 2017; Donnari et al. 2019; Akins et al. 2022; Nagaraj et al. 2022b; Kurinchi-Vendhan et al. 2023; Baes et al. 2024).
@@
Figure 14 shows the diagram of our primary galaxy sample color-coded by various dust and stellar properties. In particular, we color-coded the plane with the following physical quantities, the dust attenuation slope index (), dust optical depth in the band (), mass-weighted stellar age (), and stellar metallicity (). Instead of just showing scatter plots, which can sometimes conceal the underlying trends, we also derived the average trends of the physical properties in the diagrams, as well as their average upper uncertainty (i.e. the difference between the and percentiles of the posteriors), with the Locally Weighted Regression (LOESS) method (Cleveland & Devlin 1988) as implemented in the LOESS routine666https://pypi.org/project/loess/ by Cappellari et al. (2013). In the following figures, we use a linear local approximation, a smoothing factor of , and require that each hexbin contains at least 5 galaxies. The maps created with the LOESS method in the plane demonstrate that the effects of dust and stellar evolution can be separated.
@@
For the star-forming galaxies, we find a smooth increasing gradient of towards the upper right corner of the plane. Star-forming galaxies with bluer and colors are on average 0.13 dex younger than star-forming with redder and colors. On the other hand, we find a relatively weak trend with , with the reddest color galaxies having the lowest values. Of course, we notice that the stellar metallicity estimates for these particular galaxies have the highest uncertainties (up to 0.1 dex). We also observe that star-forming galaxies with the lowest tend to have the highest , suggesting a potential degeneracy between these two quantities. However, a strong degeneracy was not evident in the joint posterior distribution of these parameters (see discussion in Section 3.2.2).
@@
In the upper right corner of the plane, star-forming galaxies are characterized by a high dust attenuation and a relatively flat attenuation slope. This is the expected behavior since galaxies with high dust attenuation follow the attenuation curve and are red in both and colors. The flattening of the dust-attenuation curve is the result of high attenuation and a more complex geometry between the stellar and dust distributions, where more of the starlight is decoupled from dust (e.g. Charlot & Fall 2000; Narayanan et al. 2018; Salim & Narayanan 2020; Nersesian et al. 2020). Surprisingly, at the intermediate-attenuation regime (), towards the bottom left corner of the plane, galaxies seem to sit to even flatter attenuation curves. The uncertainties on the dust attenuation seem to increase with increasing dust attenuation, while the average uncertainty of the dust index remains relatively low for the star-forming galaxies ().
@@
Quiescent galaxies are relatively bluer in colors due to a very low dust attenuation that also produces very steep attenuation curves. The steepening of the dust-attenuation curve is likely a product of less dust scattering into the line of sight. This is consistent with the idea that the oldest and more passive galaxies are less dusty, and thus characterized by the lowest dust attenuation. Again, we find a smooth increasing gradient of towards the upper right corner of the plane, and also a moderate increasing trend with . Our results are consistent with earlier findings on the direction of slow aging after quenching (Whitaker et al. 2013; Leja et al. 2019c; Belli et al. 2019; Tacchella et al. 2022). In general, the oldest quiescent galaxies are also the ones with the lowest dust attenuation, steepest dust slopes, and highest . The uncertainties in the stellar and dust properties of quiescent galaxies are, on average, lower than those of star-forming galaxies. The only exception is the dust index , which exhibits high uncertainty (up to ).
@@ 4.4 The SFR100Myrs– relation
In this section we measure the relationship between SFR100Myrs and , otherwise known as the star-forming sequence (SFS; Daddi et al. 2007; Noeske et al. 2007; Elbaz et al. 2007; Leja et al. 2022). Galaxies form most of their mass on (Leitner 2012), or have passed through (Abramson et al. 2015), the SFS. Therefore, studying the SFS at different redshifts provides crucial information on the evolution of galaxies (e.g. Noeske et al. 2007; Whitaker et al. 2012, 2014; Speagle et al. 2014; Schreiber et al. 2015; Renzini & Peng 2015; Tomczak et al. 2014; Iyer et al. 2018; Leslie et al. 2020; Thorne et al. 2021; Leja et al. 2022).
@@
Overall, we find a flat relationship between SFR and for star-forming galaxies, with a weak correlation (). Quiescent galaxies form their own ‘red’ sequence, showing a stronger correlation with stellar mass (). As expected, the median SFR100Myrsvalues in bins of for star-forming galaxies (represented by open stars) align closely with the SFS defined by Leja et al. (2022), as both studies derive SFRs and stellar masses simultaneously using Prospector. However, previous studies report systematically higher SFS measurements, with offsets of approximately dex. Among these, the results from Thorne et al. (2021) show better agreement with ours, particularly at higher stellar masses (), where their SFS values are typically dex above those of Prospector. Thorne et al. (2021) also estimate SFRs and through SED fitting. In contrast, studies such as Whitaker et al. (2014), Schreiber et al. (2015), and Tomczak et al. (2016) rely on broadband photometry for their SFR and estimates. Deriving stellar properties, as done with SED fitting, tends to yield lower SFR estimates compared to traditional broadband formulae (e.g. Nersesian et al. 2019; Leja et al. 2022).
@@
Focussing on the star-forming galaxies, we find that the average trends with age and metallicity are more pronounced in the SFR100Myrs– plane, compared to what we found in the diagrams. At fixed SFR, low mass (high mass) galaxies have lower (higher) stellar metallicity, while their stellar populations are on average younger (older). Again, we find that the average uncertainties over and are relatively larger for galaxies with very low and , yet with values below 0.1 dex. Another interesting result is that the apparent – degeneracy suggested by the colors does not manifest in the SFR100Myrs– plane. In fact, the most active and massive star-forming galaxies are those with the highest dust attenuation, and with stellar metallicities very close to solar. The uncertainties on the dust attenuation for the these galaxies remain high though ().
@@
Furthermore, we find a positive trend between and the dust attenuation slope (e.g. Reddy et al. 2018; Salim et al. 2018; Barišić et al. 2020). The less massive star-forming galaxies with high SFR (i.e. high sSFR) tend to have a relatively steeper dust attenuation slope () than the more massive ones (; ). In fact, the attenuation slope seems to become shallower with decreasing sSFR among star-forming galaxies (e.g. Salim et al. 2018). Ultimately, observational (Salmon et al. 2016; Salim et al. 2018; Nersesian et al. 2020), theoretical (Witt & Gordon 2000; Charlot & Fall 2000; Calzetti 2001; Chevallard et al. 2013), and cosmological simulations (Narayanan et al. 2018; Shen et al. 2020) all point to a complex star-to-dust geometry as the main driver of the flattening of the attenuation curve.
@@
Quiescent galaxies seem to display stronger trends with both stellar and dust properties. Similarly to our findings with the diagrams, at a fixed SFR100Myrs, we find smooth increasing gradients for both stellar age and metallicity with increasing . As it has been previously shown, quiescent galaxies formed their mass early on (e.g. Tacchella et al. 2022; Kaushal et al. 2024), staying to the SFS for a very short period of time before ending up in the red sequence. At a fixed SFR100Myrs, quiescent galaxies also have steeper attenuation slopes than star-forming galaxies. At low dust optical depths, the steep dust-attenuation curve is likely the result of dust scattering, with more red light (isotropic scattering) and less blue light (forward scattering) escaping the galaxy (Chevallard et al. 2013). We find that becomes steeper with increasing stellar mass, whereas the dust optical depth is decreasing with SFR100Myrs and increasing . Indeed, we observe that the most massive quiescent galaxies are characterized by the lowest optical depths and steepest attenuation slopes.
@@ 5 Summary and conclusions
We confirm a positive trend between the light-weighted ages and the stellar velocity dispersion (). The trend with is weaker for the mass-weighted ages and . At fixed , we find hints of a dependence of on stellar age. The scaling relations presented here indicate that massive galaxies, with high tend to be older and have higher metallicity, in agreement with studies of the Local Universe (e.g. Scott et al. 2017; Li et al. 2018), reinforcing the understanding that such galaxies have undergone significant star formation and chemical enrichment during an early epoch in their history.
Quiescent galaxies are characterized by very low star formation (SFR100Myrs/) and high metallicity (). On average, quiescent galaxies in our primary sample are 1.1 Gyr older than the star-forming ones. We also find that quiescent galaxies have steeper attenuation slopes than star-forming galaxies, with the most massive quiescent galaxies being characterized by the steepest attenuation slopes. At low dust optical depths, the steep dust-attenuation curve is likely the result of dust scattering than dust absorption (e.g. Chevallard et al. 2013). We find that becomes steeper with increasing stellar mass, whereas the dust optical depth is decreasing with SFR100Myrs and increasing .
We find that low mass (high mass) star-forming galaxies have lower (higher) stellar metallicity, while their stellar populations are on average younger (older). Star-forming galaxies are also characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. We find a positive trend between and the dust attenuation slope. The less massive star-forming galaxies tend to have a relatively steeper dust attenuation slope () than the more massive ones (; ). In fact, the attenuation slope seems to become shallower with decreasing sSFR, suggesting a more complex star-to-dust geometries.
@@ Acknowledgements.
AN gratefully acknowledges the support of the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) under contract number 4000143347. AG acknowledges support from INAF-Minigrant-2022 ”LEGA-C” 1.05.12.04.01. PFW acknowledge funding through the National Science and Technology Council grant 113-2112-M-002-027-MY2. This research made use of Astropy,777http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al. 2013, 2018).References
@@
- Baes et al. (2024) Baes, M., Gebek, A., Trcka, A., et al. 2024, arXiv e-prints, arXiv:2401.04224
@@
- Gallazzi et al. (in prep.) Gallazzi, A., Zibetti, S., Kaushal, Y., et al. in prep., ApJ
@@
- Kurinchi-Vendhan et al. (2023) Kurinchi-Vendhan, S., Farcy, M., Hirschmann, M., & Valentino, F. 2023, arXiv e-prints, arXiv:2310.03083
@@
- Lewis et al. (2023) Lewis, Z. J., Andrews, B. H., Bezanson, R., et al. 2023, arXiv e-prints, arXiv:2304.12343
@@
- Woo et al. (2024) Woo, J., Walters, D., Archinuk, F., et al. 2024, arXiv e-prints, arXiv:2401.12300
@@ Appendix B Comparison with other literature measurements
In this section, we perform a direct comparison of the best-fit properties of our primary sample with those from the recent study by Cappellari (2023). Cappellari (2023) applied the PPXF method to a sub-sample of 3200 LEGA-C galaxies, using spectroscopy from the LEGA-C survey (DR3) and 28-bands photometry from two different sources, the UltraVISTA catalog by Muzzin et al. (2013a) and the COSMOS2020 catalog by Weaver et al. (2022). Cappellari (2023) used three different stellar population synthesis (SPS) models in pPXF, including the FSPS models generated with the MILES stellar library and MIST isochrones. Since there are many similarities in both our studies, in terms of input data and theoretical models, we want to investigate the consistency of the main derived properties, and whether spectroscopy is capable to constrain the stellar ages and metallicities, regardless of the fitting methods. We note here that Cappellari (2023) used a Salpeter (1955) IMF. Therefore, we rescaled the stellar masses to a Chabrier (2003) IMF by multiplying them with a constant factor of 0.61.
@@
Furthermore, we find a reasonably good agreement and small systematic offsets for the quiescent galaxies in terms of light-weighted ages (0.02 dex) and metallicities (0.07 dex). On the other hand, the ages of the star-forming galaxies that we measure are on average systematically older than those measured by Cappellari (2023), of the order of dex and with a large scatter of 0.45 dex. The offset in stellar metallicity still remains reasonable, with a mean offset of 0.17 dex, yet a large scatter of 0.38 dex. The differences in stellar ages could potentially originate from the fact that Cappellari (2023) used more photometric bands in the UV part of the SED (even including the GALEX bands), giving more weight in the blue part of the spectrum, and thus resulting in younger ages for the star-forming galaxies. Another reason could be that Cappellari (2023) uses a different definition for the stellar ages, that is the average (age), while we take the logarithm of the average age. Averaging the (age) generally yields lower values compared to taking the logarithm of the average age.

### v2

@@ 4.3 The color-color diagram
The various, complex physical processes that take place in galaxies give rise to the differences observed at UV and optical wavelengths. Interpreting the correlations between optical colors of galaxies is a challenging task. The combined effects of dust attenuation, star formation, and stellar evolution are very difficult to disentangle, leading to many degeneracies between these parameters. A simple yet powerful method to break these degeneracies is to combine an optical and an optical–NIR color. The restframe color-color diagram has proven to be quite successful in that regard (Labbé et al. 2005; Wuyts et al. 2007; Williams et al. 2009). A diagram not only can break the aforementioned degeneracies, but also can classify galaxies into quiescent and star forming. The diagram is commonly used as a quiescence diagnostic in both observed galaxy samples (e.g., Whitaker et al. 2013; Straatman et al. 2014, 2016; Barro et al. 2014; Papovich et al. 2015, 2018; Fang et al. 2018; Tan et al. 2022; Miller et al. 2022; Antwi-Danso et al. 2023; Valentino et al. 2023) and in and cosmological hydrodynamical simulations (Davé et al. 2017; Trayford et al. 2017; Donnari et al. 2019; Akins et al. 2022; Nagaraj et al. 2022b; Kurinchi-Vendhan et al. 2024; Baes et al. 2024).
@@
Figure 12 shows the diagram of our primary galaxy sample color-coded by various dust and stellar properties. In particular, we color-coded the plane with the following physical quantities, the dust attenuation slope index (), dust optical depth in the band (), mass-weighted stellar age (), and stellar metallicity (). Instead of just showing scatter plots, which can sometimes conceal the underlying trends, we also derived the average trends of the physical properties in the diagrams, as well as their average upper uncertainty (i.e., the difference between the and percentiles of the posteriors), with the Locally Weighted Regression (LOESS) method (Cleveland & Devlin 1988) as implemented in the LOESS routine888https://pypi.org/project/loess/ by Cappellari et al. (2013). In the following figures, we use a linear local approximation, a smoothing factor of , and require that each hexbin contains at least five galaxies. The maps created with the LOESS method in the plane demonstrate that the effects of dust and stellar evolution can be separated.
@@
For the star-forming galaxies, we find a smooth increasing gradient of toward the upper right corner of the plane. Star-forming galaxies with bluer and colors are on average 0.13 dex younger than star-forming with redder and colors. On the other hand, we find a relatively weak trend with , with the reddest color galaxies having the lowest values. Of course, we notice that the stellar metallicity estimates for these particular galaxies have the highest uncertainties (up to 0.1 dex). We also observe that star-forming galaxies with the lowest tend to have the highest , suggesting a potential degeneracy between these two quantities. However, a strong degeneracy was not evident in the joint posterior distribution of these parameters (see discussion in Section 3.2.2).
@@
In the upper right corner of the plane, star-forming galaxies are characterized by a high dust attenuation and a relatively flat attenuation slope. This is the expected behavior since galaxies with high dust attenuation follow the attenuation curve and are red in both and colors. The flattening of the dust-attenuation curve is the result of high attenuation and a more complex geometry between the stellar and dust distributions, where more of the starlight is decoupled from dust (e.g., Charlot & Fall 2000; Narayanan et al. 2018; Salim & Narayanan 2020; Nersesian et al. 2020). Surprisingly, at the intermediate-attenuation regime (), toward the bottom left corner of the plane, galaxies seem to sit to even flatter attenuation curves. The uncertainties on the dust attenuation seem to increase with increasing dust attenuation, while the average uncertainty of the dust index remains relatively low for the star-forming galaxies ().
@@
Quiescent galaxies are relatively bluer in colors due to a very low dust attenuation that also produces very steep attenuation curves. The steepening of the dust-attenuation curve is likely a product of less dust scattering into the line of sight. This is consistent with the idea that the oldest and more passive galaxies are less dusty, and thus characterized by the lowest dust attenuation. Again, we find a smooth increasing gradient of toward the upper right corner of the plane, and also a moderate increasing trend with . Our results are consistent with earlier findings on the direction of slow aging after quenching (Whitaker et al. 2013; Leja et al. 2019c; Belli et al. 2019; Beverage et al. 2021; Tacchella et al. 2022; Beverage et al. 2023). In general, the oldest quiescent galaxies are also the ones with the lowest dust attenuation, steepest dust slopes, and highest . The uncertainties in the stellar and dust properties of quiescent galaxies are, on average, lower than those of star-forming galaxies. The only exception is the dust index , which exhibits high uncertainty (up to ).
@@ 4.4 The SFR100Myrs– relation
In this section we measure the relationship between SFR100Myrs and , otherwise known as the star-forming sequence (SFS; Daddi et al. 2007; Noeske et al. 2007; Elbaz et al. 2007; Leja et al. 2022). Galaxies form most of their mass on (Leitner 2012), or have passed through (Abramson et al. 2015), the SFS. Therefore, studying the SFS at different redshifts provides crucial information on the evolution of galaxies (e.g., Noeske et al. 2007; Whitaker et al. 2012, 2014; Speagle et al. 2014; Schreiber et al. 2015; Renzini & Peng 2015; Tomczak et al. 2014; Iyer et al. 2018; Leslie et al. 2020; Thorne et al. 2021; Leja et al. 2022).
@@
Overall, we find a flat relationship between SFR and for star-forming galaxies, with a weak correlation (). Quiescent galaxies form their own ‘red’ sequence, showing a stronger correlation with stellar mass (). As expected, the median SFR100Myrsvalues in bins of for star-forming galaxies (represented by open stars) align closely with the SFS defined by Leja et al. (2022), as both studies derive SFRs and stellar masses simultaneously using Prospector. However, previous studies report systematically higher SFS measurements, with offsets of approximately dex. Among these, the results from Thorne et al. (2021) show better agreement with ours, particularly at higher stellar masses (), where their SFS values are typically dex above those of Prospector. Thorne et al. (2021) also estimate SFRs and through SED fitting. In contrast, studies such as Whitaker et al. (2014), Schreiber et al. (2015), and Tomczak et al. (2016) rely on broadband photometry for their SFR and estimates. Deriving stellar properties, as done with SED fitting, tends to yield lower SFR estimates compared to traditional broadband formulae (e.g., Nersesian et al. 2019; Leja et al. 2022).
@@
Focussing on the star-forming galaxies, we find that the average trends with age and metallicity are more pronounced in the SFR100Myrs– plane, compared to what we found in the diagrams. At fixed SFR, low-mass (high-mass) galaxies have lower (higher) stellar metallicity, while their stellar populations are on average younger (older). Again, we find that the average uncertainties over and are relatively larger for galaxies with very low and , yet with values below 0.1 dex. Another interesting result is that the apparent – degeneracy suggested by the colors does not manifest in the SFR100Myrs– plane. In fact, the most active and massive star-forming galaxies are those with the highest dust attenuation, and with stellar metallicities very close to solar. The uncertainties on the dust attenuation for these galaxies remain high though ().
@@
Furthermore, we find a positive trend between and the dust attenuation slope (e.g., Reddy et al. 2018; Salim et al. 2018; Barišić et al. 2020). The less massive star-forming galaxies with high SFR (i.e., high sSFR) tend to have a relatively steeper dust attenuation slope () than the more massive ones (; ). In fact, the attenuation slope seems to become shallower with decreasing sSFR among star-forming galaxies (e.g., Salim et al. 2018). Ultimately, observational (Salmon et al. 2016; Salim et al. 2018; Nersesian et al. 2020), theoretical (Witt & Gordon 2000; Charlot & Fall 2000; Calzetti 2001; Chevallard et al. 2013), and cosmological simulations (Narayanan et al. 2018; Shen et al. 2020) all point to a complex star-to-dust geometry as the main driver of the flattening of the attenuation curve.
@@
Quiescent galaxies seem to display stronger trends with both stellar and dust properties. Similarly to our findings with the diagrams, at a fixed SFR100Myrs, we find smooth increasing gradients for both stellar age and metallicity with increasing . As it has been previously shown, quiescent galaxies formed their mass early on (e.g., Tacchella et al. 2022; Kaushal et al. 2024), staying to the SFS for a very short period of time before ending up in the red sequence. At a fixed SFR100Myrs, quiescent galaxies also have steeper attenuation slopes than star-forming galaxies. At low dust optical depths, the steep dust-attenuation curve is likely the result of dust scattering, with more red light (isotropic scattering) and less blue light (forward scattering) escaping the galaxy (Chevallard et al. 2013). We find that becomes steeper with increasing stellar mass, whereas the dust optical depth is decreasing with SFR100Myrs and increasing . Indeed, we observe that the most massive quiescent galaxies are characterized by the lowest optical depths and steepest attenuation slopes.
@@ 5 Summary and conclusions
We confirm a positive trend between the light-weighted ages and the stellar velocity dispersion (). The trend with is weaker for the mass-weighted ages and . At fixed , we find hints of a dependence of on stellar age. The scaling relations presented here indicate that massive galaxies, with high tend to be older and have higher metallicity, in agreement with studies of the Local Universe (e.g., Scott et al. 2017; Li et al. 2018), reinforcing the understanding that such galaxies have undergone significant star formation and chemical enrichment during an early epoch in their history.
Quiescent galaxies are characterized by very low star formation (SFR100Myrs/) and high metallicity (). On average, quiescent galaxies in our primary sample are 1.1 Gyr older than the star-forming ones. We also find that quiescent galaxies have steeper attenuation slopes than star-forming galaxies, with the most massive quiescent galaxies being characterized by the steepest attenuation slopes. At low dust optical depths, the steep dust-attenuation curve is likely the result of dust scattering than dust absorption (e.g., Chevallard et al. 2013). We find that becomes steeper with increasing stellar mass, whereas the dust optical depth is decreasing with SFR100Myrs and increasing .
We find that low-mass (high-mass) star-forming galaxies have lower (higher) stellar metallicity, while their stellar populations are on average younger (older). Star-forming galaxies are also characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. We find a positive trend between and the dust attenuation slope. The less massive star-forming galaxies tend to have a relatively steeper dust attenuation slope () than the more massive ones (; ). In fact, the attenuation slope seems to become shallower with decreasing sSFR, suggesting a more complex star-to-dust geometries.
@@ Acknowledgements.
We thank the anonymous referee for the valuable remarks and suggestions that helped us improve the paper. AN gratefully acknowledges the support of the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) under contract number 4000143347. AG acknowledges support from INAF-Minigrant-2022 ”LEGA-C” 1.05.12.04.01. PFW acknowledge funding through the National Science and Technology Council grant 113-2112-M-002-027-MY2. This research made use of Astropy,999http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al. 2013, 2018).References
@@
- Baes et al. (2024) Baes, M., Gebek, A., Trčka, A., et al. 2024, A&A, 683, A181
@@
- Beverage et al. (2021) Beverage, A. G., Kriek, M., Conroy, C., et al. 2021, ApJ, 917, L1
- Beverage et al. (2023) Beverage, A. G., Kriek, M., Conroy, C., et al. 2023, ApJ, 948, 140
@@
- Kurinchi-Vendhan et al. (2024) Kurinchi-Vendhan, S., Farcy, M., Hirschmann, M., & Valentino, F. 2024, MNRAS, 534, 3974
@@
- Lewis et al. (2024) Lewis, Z. J., Andrews, B. H., Bezanson, R., et al. 2024, ApJ, 964, 59
@@
- Woo et al. (2024) Woo, J., Walters, D., Archinuk, F., et al. 2024, MNRAS, 530, 4260
@@ Appendix B Comparison with other literature measurements
In this section, we perform a direct comparison of the best-fit properties of our primary sample with those from the recent study by Cappellari (2023). Cappellari (2023) applied the PPXF method to a subsample of 3200 LEGA-C galaxies, using spectroscopy from the LEGA-C survey (DR3) and 28-bands photometry from two different sources, the UltraVISTA catalog by Muzzin et al. (2013a) and the COSMOS2020 catalog by Weaver et al. (2022). Cappellari (2023) used three different stellar population synthesis (SPS) models in pPXF, including the FSPS models generated with the MILES stellar library and MIST isochrones. Since there are many similarities in both our studies, in terms of input data and theoretical models, we want to investigate the consistency of the main derived properties, and whether spectroscopy is capable to constrain the stellar ages and metallicities, regardless of the fitting methods. We note here that Cappellari (2023) used a Salpeter (1955) IMF. Therefore, we rescaled the stellar masses to a Chabrier (2003) IMF by multiplying them with a constant factor of 0.61.
@@
Furthermore, we find a reasonably good agreement and small systematic offsets for the quiescent galaxies in terms of light-weighted ages (0.02 dex) and metallicities (0.07 dex). On the other hand, the ages of the star-forming galaxies that we measure are on average systematically older than those measured by Cappellari (2023), on the order of dex and with a large scatter of 0.45 dex. The offset in stellar metallicity still remains reasonable, with a mean offset of 0.17 dex, yet a large scatter of 0.38 dex. The differences in stellar ages could potentially originate from the fact that Cappellari (2023) used more photometric bands in the UV part of the SED (even including the GALEX bands), giving more weight in the blue part of the spectrum, and thus resulting in younger ages for the star-forming galaxies. Another reason could be that Cappellari (2023) uses a different definition for the stellar ages, that is the average (age), while we take the logarithm of the average age. Averaging the (age) generally yields lower values compared to taking the logarithm of the average age.
@@
Appendix C Example of joint posterior distributions for a star-forming galaxy
In this section, we present the resulting joint posterior distributions of the main physical quantities (Fig. 16), and a comparison of the posteriors retrieved from three different SED fitting runs (Fig. 17), of a handpicked star-forming galaxy with the LEGA-C ID 2808. See main text for the discussion.